# 2D Entanglement from a Single Motzkin Spaghetto

### v1

@@ Abstract
TN—The GS of the 1D Motzkin chain has a holographic TN representation [24], where each physical spin maps to a column of tensors sequentially contracted with neighbors above and below. The depth of the columns grows linearly from the boundaries of the chain to the center. Twisting the TN from the center results in a 2D TN for the GS of the present model, which has a pyramid shape. The geometry of the TN is globally the same as the 3D TN for the GS of the coupled Fredkin chain model [26]. But the tensors in the bulk have only four internal legs instead of six. The number of bonds crossing the bipartition cut gives the correct scaling of the EE for the $q>1$ phase of the colored model. At the critical point of the colorless model, the TN can be renormalized such that the depth of the columns grow logarithmically instead, giving the correct EE scaling in that case [25].
@@
Conclusion—IN this letter, a frustration free 2D lattice Hamiltonian with exactly solvable GS that has volume scaling of EE is constructed by twisting the highly entangled Motzkin chain. The interaction is significantly simpler than the previous and only model of such highly entangled 2D systems realized by coupling arrays of highly entangled chains. Therefore it has relatively better prospect in experimental realization and applications in quantum control and simulation. A spin-$1/2$ version of the model can be constructed by twisting a Fredkin completely analogously. At the moment, it is not clear if similar construction works for 3D systems, whereas the coupled chains construction in principle works in any dimension via tiling. These alternative examples of area-law violating GSs provide a better understanding of the interplay between EE and spectral gap scaling, or the tradeoff between quantum computing resource and robustness against noise. The ultimate goal in this direction is to generalize the area law to gapless ground states to get a relation between the rate of decay in spectral gap and growth of EE with the system size.
@@ Acknowledgements.
I thank Shankar Balasubramanian, Soonwan Choi, Israel Klich, Ethan Lake and Olai Mykland for discussions.

### v2

@@ IV Antiferromagnetic order
The GS of the 1D Motzkin chain has a holographic TN representation [35], where each physical spin maps to a column of tensors sequentially contracted with neighbors above and below. The depth of the columns grows linearly from the boundaries of the chain to the center. Twisting the TN from the center results in a 2D TN for the GS of the present model, which has a pyramid shape as shown in Fig. 3 (a). The geometry of the TN is globally the same as the 3D TN for the GS of the coupled Fredkin chain model [29] shown in Fig. 3 (b). But now the tensors in the bulk have only four internal legs instead of six. The number of bonds crossing the bipartition cut gives the correct scaling of the EE for the $q>1$ phase of the colored model. At the critical point of the colorless model, the TN can be renormalized such that the depth of the columns grow logarithmically instead, giving the correct EE scaling in that case [26, 27].
The TN representation of the Motzkin and Fredkin GS defined in [35] actually gives an inverted step pyramid shaped network with the physical external legs at the bottom of each column of tensors. In this section, we define a new TN representation of the Fredkin GS with an upright pyramid shaped geometry, which consist of slightly sparser tensors. This TN is obtained using the tiling approach used in Ref. [35]. In this method the correspondence between spin configurations and walks is used to construct the TN representation. This is achieved through an intermediate step where a valid tiling of a set of tiles is associated to each walk present in the GS. In this context, a valid tiling is a tiling where each edge of each tile matches its adjacent tiles. By carefully choosing an appropriate set of tiles, one can assure a one-to-one correspondence between valid tilings and spin configurations present in the GS superposition. For spin configurations present in the GS of the Fredkin model, which are described by Dyck walks, this one-to-one correspondence can be assured by the set of tiles $B_{i}(\bm{c})$ seen in panel (a) of figure 4. In panel (b) and (c) we see that the valid tilings contain arrowed paths indicating the color correlation between spins in the given Dyck walk. A colored spin up (resp. down) at $i$ is represented by a colored arrowed path entering (resp. leaving) the tiling at the floor edge of the bottom tile in the tower of tiles at $i$. Clearly, each valid tiling corresponds to a unique way of pairing the degrees of freedom. This set of tiles enables exactly one valid tiling for any given Dyck walk and no valid tiling for non-Dyck walks. To see this, note that by following the highest arrowed path within each tiling, the corresponding Dyck walk is traced out, see figure 4 (b) and (c). It is clear that a valid tiling for a given Dyck walk where this is not the case, is not possible. It is also clear that no valid tiling can be constructed for a non-Dyck walk, as that would necessarily violate arrow continuity at the step where the walk becomes negative (or at the endpoints if the walk starts or terminates at a non-zero height). This establishes the one-to-one correspondence between valid tilings and the spin configurations in the ground state.
The tensor building up the TN representation is then defined by the set of tiles in the following way: The number of indices of the tensor is determined by the shape of the tiles and each tile represents an index configuration of the tensor giving a nonzero value. Contractions between two tensors in the TN, then correspond to summing over tile configurations matching at that edge. The tiles, valid tilings and the resulting TN, can be seen in figure 4.
[…]
Figure 4: (a) The five different tiles $B_{i}(\bm{c})$ for the new Fredkin tensor network. The tiles are also defined as rank-4 tensors in terms of Kronecker deltas, where the indices $\bm{k}_{i}$ are defined as for $B(q)$ in panel (d). We have $\bm{c}=(1,0)$ for red arrow, $\bm{c}=(0,1)$ for blue arrow. No arrow corresponds to $\bm{0}$. (b)-(c) Valid tilings corresponding to the maximal and minimal height Dyck walk for the $L=4$ system. (d) TN representation of the ground state of the single Fredkin chain for $L=4$. The constituent tensors are defined in the upper panel. $l$ denotes the different levels of the holographic TN and the white circle projects out the index value $\bm{0}$ corresponding to no arrow in the tilings.
The nonzero entries of the tensor in terms of the deformation parameter $q$ are determined by counting the number of tiles in a tiling and considering the area under the corresponding Dyck walk. By weighing each tile with an appropriate value of the deformation parameter $q$, one can ensure that contracting the TN yields the weighted superposition of spin configurations in the GS. This gives a rank-4 tensor $B(q)$ as
$\begin{split}B(q)=\sum_{\bm{c}_{i}=(1,0),(0,1)}&[\sqrt{q}B_{1}(\bm{c}_{1})+qB_{2}(\bm{c}_{1},\bm{c}_{2})\\ +&qB_{3}(\bm{c}_{1},\bm{c}_{2}),+\sqrt{q}B_{4}(\bm{c}_{1})+B_{5}(\bm{c}_{1})].\end{split}$ (5)
The indices $\bm{k}_{1},\cdots,\bm{k}_{4}$ of $A(q)$ are defined in panel (d) of figure 4 and have been suppressed in equation 5. We use a vector notation for the indices to keep track of the color of the arrows, which is a notation conveniently extended to the tensors in the TN for the 2D systems [29]. It is clear that one can simply replace the different vectors with a unique number, to arrive at the more familiar way of denoting index configurations by numbers. Note also that we use a minus sign to indicate arrows leaving (resp. coming in) at the $\bm{k}_{1}$ (resp. $\bm{k}_{3}$) index. In equation 5, we consider two colors, namely red, $\bm{c}=(1,0)$, and blue, $\bm{c}=(0,1)$. Note that each of the tiles $B_{i}(\bm{c})$ can be defined as a rank-4 tensor expressed as a product of Kronecker deltas, as seen in panel (a) in figure 4. The shape of the resulting TN naturally becomes that of the valid tilings, giving an upright step pyramid shape as seen in panel (d) in figure 4. Note that the boundary tensors $\delta_{\bm{k},\bm{0}}$ ensures that “no arrows flow out” of the TN, except at the bottom edges. Due to the tile without arrows, we must project out the index value corresponding to an empty edge at the bottom tile. We must therefore include projectors in the TN, as seen in figure 4 (d). This is similar to the rainbow TN presented with projectors in Ref. [35].
Finally, we mention that one can restrict the tensors at even and odd sites by taking into account that a Dyck walk can only reach an even (resp. odd) height after an even (resp. odd) number of steps. This implies that in valid tilings the tile above an even (resp. odd) position $i$, at an odd (resp. even) level $l$ can never be the $B_{1}(\bm{c})$ tile, as that would correspond to a walk reaching an even height (resp. odd) after an odd (resp. even) number of steps. Due to the same reason the $B_{4}(\bm{c})$ tile can never be placed above an even (resp. odd) position, at an even (resp. odd) height. This means that one can build a TN where the tensors are built up of 4 tiles, instead of 5, at the expense of having two different tensors. One is $B(q)$ without $B_{1}(\bm{c})$ and the other one is $B(q)$ without $B_{4}(\bm{c})$. This gives a TN with slightly sparser tensor than the one presented in Ref. [35], at the cost of having two different tensors in the TN.
VI Conclusion
@@
In this article, a frustration free 2D lattice Hamiltonian with exactly solvable GS that has volume scaling of EE is constructed by twisting the highly entangled Motzkin chain. The interaction is significantly simpler than the previous and only model of such highly entangled 2D systems realized by coupling arrays of highly entangled chains. Similar constructions of 2D entanglement from twisting 1D chains can be applied to the highly entangled XX chain with inhomogeneous coupling [8, 9, 10]. At the moment, it is not clear if similar construction works for 3D systems, whereas the coupled chains construction in principle works in any dimension via tiling. These alternative examples of area-law violating GSs provide a better understanding of the interplay between EE and spectral gap scaling, or the tradeoff between quantum computing resource and robustness against noise. The ultimate goal in this direction is to generalize the area law to gapless ground states to get a relation between the rate of decay in spectral gap and growth of EE with the system size.
@@ Acknowledgements.
ZZ thanks Shankar Balasubramanian, Soonwan Choi, Israel Klich, and Ethan Lake for discussions.